OpenStudy (anonymous):

Find (fog)(2) and (f+g)(2) when f(x)=1/x and g(x)=4x+9.

OpenStudy (anonymous):

(fog)(2)=f[g(2)] =f(17) =1/17

OpenStudy (anonymous):

(f+g)(2) = f(2) + g(2) =(1/2)+17 =17.5